OpenStudy (anonymous):

Which polynomial represents a difference of cubes? x^3 -64 4x^3 -1 8x^3 +27 4x^3 -125

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

a cube is the result when a number or variable is multiplied by itself 3 times for example the cube of x = x*x*x = x^3 the cube of 2 is 8 because 2 * 2* 2 = 8 and 2 is called the cube root of 8

OpenStudy (welshfella):

a difference, of course, is when one number or variable is subtracted from another

OpenStudy (welshfella):

So which one do you think is difference of 2 cubes?

OpenStudy (welshfella):

a few more cubes are 3^3 = 27 4^3 = 64 5^3 = 125

OpenStudy (anonymous):

x^3-64?

OpenStudy (welshfella):

right

OpenStudy (welshfella):

is 4x^3 a perfect cube?

OpenStudy (anonymous):

yea

OpenStudy (welshfella):

Ah caught you there! x^3 is one but 4 is not 4 is a perfect square But 8x^3 is a cube 2x*2x*2x = 8x^3

OpenStudy (anonymous):

lol oh ok. Thank u.

OpenStudy (welshfella):

yw